OpenStudy (anonymous):

Halley drank 2 1/3 glasses of milk on Monday. Tyrone drank 1 1/3 times as much milk as Halley. How many glasses of milk did Tyrone drink?

OpenStudy (anonymous):

(About. NOT exactly) 24.7 oz. So, Tyrone drank (about) 3 cups of milk. 3-2 1/3= 2/3.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Yep, 3 sounds accurate.

OpenStudy (phi):

you should multiply the two numbers 2 and 1/3 times 1 and 1/3 first write them as an "improper fraction" 2 and 1/3 is 7/3 (do you know how to do that ?) and 1 1/3 is 4/3 now do \[ \frac{7}{3}\cdot \frac{4}{3} \] multiply top times top and bottom times bottom to get \[ \frac{28}{9}\]